OpenStudy (anonymous):

Find an exact value. cos 15°

OpenStudy (arindameducationusc):

cos(45-30) using cos(A-B)=cosAcosB+sinAsinB

OpenStudy (arindameducationusc):

Got?

OpenStudy (anonymous):

cos (45 - 30) = cos45cos30 + sin45sin30 ?

OpenStudy (arindameducationusc):

yes

OpenStudy (arindameducationusc):

Can you solve now?

OpenStudy (anonymous):

so then would it be cos (45-30) = (radical) 2 / 2 * (radical) 3 / 2 + (radical) 2 / 2 * 1 / 2 ?

OpenStudy (anonymous):

would it be \[\sqrt{6} + \sqrt{2} / 4\]

OpenStudy (arindameducationusc):

It should be Sqrt[3] - 1/2 Sqrt[2]

OpenStudy (arindameducationusc):

1/sqrt(2)*sqrt(3)/2 - 1/sqrt(2)*1/2

OpenStudy (arindameducationusc):

=sqrt(3)-1 / 2sqrt(2)

OpenStudy (arindameducationusc):

Got it?

OpenStudy (anonymous):

thats not one of my answer choices :(

OpenStudy (anonymous):

these are my answer choices

OpenStudy (arindameducationusc):

okay let me see....

OpenStudy (anonymous):

@lxoser cos(15) To find this value, all you have to do is express 15 as a sum or difference of two known unit circle values. Then, take advantage of the sum/difference identities. cos(a + b) = cos(a)cos(b) - sin(a)sin(b) cos(a - b) = cos(a)cos(b) + sin(a)sin(b) [note: watch how the addition one uses subtraction, and subtraction uses addition] For this case, 15 = 45 - 30. Therefore, cos(15) = cos(45 - 30) And we know what the sine and cosine of 45 degrees and 30 degrees are. Using the cosine difference identity, we get cos(45)cos(30) + sin(45)sin(30) Which we know how to solve. (sqrt(2)/2) (sqrt(3)/2) + (sqrt(2)/2) (1/2) Keep simplifying, [ sqrt(6)/4 ] + [ sqrt(2) / 4 ] [sqrt(6) + sqrt(2)] / 4

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

so i got it right the first time, thank you so much guys for your time & help! i understand now :-)

OpenStudy (anonymous):

i live in florida

OpenStudy (anonymous):

is that FLVS?

OpenStudy (anonymous):

lol, yes it is

OpenStudy (anonymous):

i completely agree!! lol

OpenStudy (anonymous):

i like how if you miss one you get an 80 on the homework

OpenStudy (anonymous):

yeah i know! like really?

OpenStudy (anonymous):

and it brings your grade down like 2-3 points but if you get a 100 it brings it up like .25

OpenStudy (anonymous):

lmao omg i feel your pain so much

OpenStudy (anonymous):

what math are you in

OpenStudy (anonymous):

what math are you in

OpenStudy (anonymous):

im in pre-cal honors

OpenStudy (anonymous):

if you go to google and type in flvs it got a 2.5 on google but flvs only shows the good ones on the side XD

OpenStudy (anonymous):

lol, really?? florida students got it rough lol